# Johannes Huemer [^2] These authors contributed equally to this work. Markus Murschitz33footnotemark: 3 Matthias Schörghuber Lukas Reisinger Thomas Kadiofsky Christoph Weidinger Mario Niedermeyer Benedikt Widy Marcel Zeilinger Csaba Beleznai Tobias Glück Andreas Kugi Patrik Zips44footnotemark: 4

### v1

@@ 3.2 Actuation
As shown in Table 1, the BM154 platform uses a diesel engine as its main power source, driving two distinct hydraulic circuits that control vehicle movement and operational functions. The diesel engine drives a closed hydraulic circuit designed for operating the wheels, as well as an open hydraulic circuit responsible for actuating multiple hydraulic cylinders, amongst others for steering and lifting the forks. Figure 4 and Table 2 provide graphical and concise textual information on the forklift’s actuated and non-actuated joints, driven by the hydraulic motor and cylinders.
The closed hydraulic circuit is tasked with driving the vehicle’s wheels, delivering continuous hydraulic traction to ensure efficient power transfer and enhanced mobility. This design optimizes vehicle stability and maneuverability, particularly in challenging or uneven terrains. The hydrostatic four-wheel drive system is controlled by a PWM signal, where the duty cycle determines the flow rate of the hydraulic fluid, which in turn directly governs the wheel speed.
@@ 3.3.1 Proprioception
The forklift’s proprioception system integrates several sensors to monitor its internal state with high precision, especially the joint values depicted in Figure 4. In terms of vehicle motion feedback, wheel speed encoders are used to monitor the rotational velocity of each wheel. To capture the 3D rotation of the central articulated joint, a system utilizing three linear potentiometers was designed. By measuring the sensor elongation, the three articulation angles can be calculated with sub-degree precision. The resulting values are integral for steering the vehicle and precisely adjusting the fork tilt angle during load manipulation. Pressure sensors are installed within the hydraulic cylinders and the hydraulic drive system to measure pressure levels at high frequency. These measurements are crucial for controlling the forklift’s load-handling capabilities, including detecting whether the fork is in direct contact with the environment. In addition, draw-wire encoders are implemented to measure the vertical position of the lifting mast and the lateral movement of the forks (side-shift) with millimeter precision. Together, these sensors provide comprehensive real-time feedback and allow for accurate modeling of the behavior of the system.
@@ 3.3.2 Exteroception
For pallet recognition, the ZED2i is the central component. It provides the input data for the neural networks for pallet recognition and pose estimation. Additionally, for precise short-range detection during the final pallet approach, the forklift is equipped with a Sick picoScan150 2D LiDAR, Figure. 3d. It ensures accurate positioning of the forks relative to the pallet, due to its high frequent update rate and well-aligned viewing angle. To enable accurate detection of the truck’s loading edge, the forklift utilizes the Livox LiDAR. Additionally, a Septentrio mosaic-H system with two antennas is used for RTK GNSS localization. It provides accurate position information at the centimeter level and heading information to ensure precise alignment during loading and unloading tasks.
@@ 3.4 Processing Components and Network
The B&R X90 PLC is dedicated to managing low-level hardware interfaces, including direct connections to sensors and actuators that require low latency and real-time control. It also handles safety-critical operations that depend on rapid response times, such as emergency stops. The rugged IP67-compliant design of the PLC makes it suitable for harsh outdoor environments and does not require an additional enclosure. For higher-level processing, such as planning, advanced control algorithms, and managing intelligent sensors, a Vecow 1210 IPC is employed. It connects to the PLC via a central Ethernet switch to receive the low-level sensor data and to transmit high-level commands over a custom TCP/IP protocol. This setup enables seamless integration and coordination between low-level operations and advanced computational tasks that do not require millisecond-level latency. These advanced computational tasks include high-level task and motion control, vehicle state estimation, and object detection pipelines, utilizing camera and LiDAR sensors. The NUC unit is responsible for processing data from the Ouster LiDAR, allowing both mapping the vehicle’s environment and obstacle avoidance and sharing the respective data via the Robot Operating System (ROS 2) [65] over Ethernet.
A wireless network bridge connects the forklift to a remote monitoring station housed in a weatherproof container, primarily supporting development and expert interventions. This setup enables live monitoring of the system’s operation, ensuring efficient debugging and adjustments during testing phases. Additionally, a supplementary WLAN antenna provides on-site operators with the ability to monitor the system state via a tablet-based Human-Machine Interface (HMI). For long-range communication, an LTE/5G router is integrated.
@@ 3.5 Software Components
Figure 2 shows an overview of the software architecture as well as concise descriptions of core functionalities and their associated hardware components. Core functionalities include localization and mapping, object detection, obstacle detection, task planning, and motion control. Localization and mapping are based on GNSS modules and LiDAR to provide accurate environmental maps and vehicle positioning data. Object detection utilizes a stereo camera as well as LiDAR to identify pallets and truck loading edges, facilitating precise load carrier handling. The obstacle detection module ensures safety by halting the vehicle before potential collisions with detected obstacles. The task planner serves as the central planning component of the system, managing high-level decision-making and coordinating both the path planner and motion controller. The path planner creates optimal routes based on map and localization data, while the cascaded motion controller provides precise vehicle base movements and responsive fork actuation for effective pallet manipulation. The majority of the software components running on the rugged IPC and the NUC highly depend on the ROS 2 middleware and its associated libraries, which enable a decentralized architecture while ensuring robust data and information exchange. A detailed discussion of the modules presented can be found in Sections 4 and 5.
@@ 4.1 Joint Localization and Pallet Mapping
The goal of mapping is to derive a representation of the environment that is as rich in features as necessary for the application, but as compact as possible to keep computational complexity low. For ADAPT, there are two levels of abstraction in which we need to map the environment in: The high-level representation of the pallet poses, and the lower-level environment map with a strong focus on terrain traversability assessment, see Section 4.2.
@@
For high-level joint location and pallet mapping, we use a factor graph-based approach [66] that allows the integration of vehicle odometry data, GNSS measurements and pallet detections in a joint optimization framework. A factor graph is a theoretical framework that models unknown variables, measurements, and their interdependencies as a bipartite graph with two types of nodes, see Figure 6. It consists of variable nodes (depicted as circles), which represent the variables to be estimated, and factor nodes (depicted as rectangles). In this context, factors can be binary (functions that relate two unknowns with one another) or unary (constraints on a variable based on a measurement). This factor graph approach is advantageous in this context as it allows flexible integration of multiple types of measurement and provides a probabilistic framework for managing uncertainties inherent in noisy and delayed sensor data. Our factor graph is shown in Figure 6. We use GNSS measurements from a dual antenna system that provides global position and attitude information. The forklift pose nodes are connected using odometry which is based on wheel encoders and steering angle measurements. Pallets are detected and their poses are estimated from different viewpoints and forklift poses, respectively. Therefore, the estimation of the pose of the pallet is represented as a binary factor linking the position of the pallet with the location of the vehicle. To maintain real-time performance, we build upon iSAM2 [33] incremental smoothing, which yields the jointly most probable pallet and forklift poses given all odometry, GNSS, and pallet measurements. This allows the pose of a pallet to be optimized by measurements from different views. In addition, the pallet poses and vehicle location are optimized even without or with a less precise GNSS signal.
@@
With iSAM2 updates of pallet detections and localization information are optimized incrementally, in contrast to re-solving the entire optimization repeatedly. To do so, iSAM2 maintains, in addition to the factor graph, a dynamic Bayesian tree structure to track variable interdependencies. This approach ensures that updates to the factor graph are confined to regions where changes in interdependencies occur. To facilitate real-time operation, our iSAM2-based optimization approach for joint pallet and vehicle pose mapping
@@
  • exploits the sparsity of the optimization problem, since each factor connects only a small subset of variables, leading to a sparse Jacobian matrix,
@@
This approach can efficiently integrate information from odometry sensors, GNSS, and detections over time in a way that yields, given all sensor information, the most probable result for the unknowns (the maximum a posteriori estimate). The pallet mapping is tightly integrated into our optimization framework. For pallet tracking, we use a Mahalanobis distance-based [67] approach to ensure a robust association of detections between consecutive measurements. The Mahalanobis distance metric allows for dynamic association by accounting for both positional variance and estimated noise in pallet positions. Additionally, pallets are dynamically managed within the forklift’s view frustum by pallet bookkeeping that goes beyond pure variable marginalization (pruning). It manages the tracking of new never-before-seen pallets, associates new detections with old ones, and deletes pallets if they should be in sensor view but cannot be detected for a while. This approach aligns with the concept of object permanence, where pallets are presumed to remain at a designated location even when they are not directly visible.
@@
Figure 6: Factor Graph for the joint forklift poses ($X_{x}$) and pallet poses ($P_{p}$) over time (left to right), relative (from one time step to the next) odometry information $O_{p1,p2}$, and GNSS measurements $G_{g}$. In this example, the pallet $P_{0}$ has been detected in timestep $1$ and $2$, in timestep 00 and $2$ GNSS information was available.
4.2 Traversabilty Mapping
@@
The targeted area of application, e.g., construction sites, of ADAPT are unstructured environments with potentially uneven terrain. Hence, a specific environment mapping targeting these aspects is required. Therefore, a 2.5D elevation map representation is generated from the 3D point measurements of the Ouster LiDAR. While path planning complexity can be reduced to planning in 2 dimensions, the most relevant geometric information about the environment must still be available from the map. In particular, it must allow for a clear distinction between traversable areas, obstacles, and slopes. This distinction is crucial for forklifts, as they are more susceptible to tipping over when carrying loads.
Similarly to spatial-temporal voxel layers [20], our approach aggregates data into a 3D voxel grid with short-term memory. However, instead of directly deriving a 2D costmap, our method employs a 2.5D elevation map for dynamic obstacle handling and terrain analysis, serving as a long-term memory representation. Therefore, it is implemented as a three-stage process, encompassing the analysis of single scans, 3D voxel map aggregation, and the generation of a 2.5D elevation map. In all three steps, distinguishing between ground points and obstacle points plays a decisive role.
Analysis of Single Scans: Single scans are analyzed within the sensor domain, utilizing the unique capabilities and characteristics of the sensor. By focusing on individual scans, this stage minimizes disturbances caused by alignment errors. Parameters such as the height above ground and the angle between the surface normal vector and the vertical axis help determine the probability of a point being an obstacle. For the Ouster LiDAR, a single swipe of the nearly 180° field of view that is used constitutes a scan. The ego-motion during this rotation is compensated by transforming all points into the sensor’s coordinate system at a fixed timestamp.
@@ 4.3 Pallet Detection and Pose Estimation
Detecting pallets and accurately determining their 6D pose is essential to enable ADAPT to interact with them effectively. This is particularly important in construction sites, where, unlike in controlled environments such as warehouses, pallets are typically placed arbitrarily rather than following a fixed grid or placement schedule.
@@
Pallet representation: Pallets are represented by a part-constellation model, see Figure 9(a), where pallet corners and edges not only form a localized part, but at every pixel they also define a spatial vote offset attribute, implying the offset vector to the corresponding object center. Part locations and local vote vectors are learned from synthetic data.
@@
Data generation: A synthetic data generation pipeline is used where more than 150 thousand stereo image pairs are generated with diverse pallet configurations, e.g., orientations, items carried, clutter. The data generation step relies on a simulated stereo camera setup, matching the baseline and optics of a ZED2i stereo camera [68]. The synthetic stereo image pairs are used in a stereo depth computation step, which results in depth data exhibiting similar data characteristics as depth computed from real stereo pairs. This low sim-to-real gap allows for generating a rich variety of view variations and pallet configurations for learning.
Learning: We employ an extensible convolutional encoder-decoder framework, based on [69], which estimates outputs for multiple learning tasks, see Figure 8. Pallets are represented by a part-based representation, where each part also carries a spatial vote offset attribute, where the offset vector points to the corresponding object center. During training key-point and line segment representations yield multichannel heatmap output representations, where each channel is assigned to respective corner/edge categories.
@@
Pose estimation accuracy: By analyzing the geometry of Euro-pallets and fork dimensions, precise tolerance values for pallet insertion can be established. Consequently, pose estimation errors serve as a key metric for evaluating the effectiveness of the proposed approach in achieving accurate insertion. The ground truth data for error computation was obtained through manual annotation of a point cloud captured by an extrinsically calibrated LiDAR sensor mounted adjacent to the camera. Figure 10 presents the evaluation results based on 170 detections, covering various load types at different distances and orientations. To compare a detection with its corresponding annotation, the estimated 6D pose was transformed into a coordinate system centered at the manually annotated pose, cf. Figure 9(b), yielding a 6D error vector $\mathbf{e}=[e_{x},e_{y},e_{z},e_{\phi},e_{\theta},e_{\psi}]^{\mathrm{T}}$. The tolerance limits for the relevant degrees of freedom are $y_{tol}=\pm 0.05m$ and $z_{tol}=\pm 0.04m$, with respect to the pallet coordinate system. The constraint in the x-direction is theoretically determined by the load’s center of mass, however, complete fork insertion is preferable to ensure optimal transport stability. Assuming the pallets remain parallel to the ground, detections with unsuitable roll and pitch angles are filtered out, as the forklift cannot successfully pick them up. Thus, these dimensions are excluded from the evaluation. Errors in the y-direction $e_{y}$ are affected by deviations in orientation around the pallet’s z-axis, which modify the effective pallet opening by a factor of $\cos(e_{\psi})$. To account for this influence, the corrected y-direction error is computed as $e_{y}^{\psi}=\frac{e_{y}}{\cos(e_{\psi})}$.
The results indicate that detection accuracy deteriorates with increasing sensor distance, particularly in $e_{x}$, due to the quadratic increase in depth error inherent in stereo vision. Furthermore, when the pallet is rotated around its z-axis, the error distributes between $e_{x}$ and $e_{y}$. Overall, the results suggest that a single detection at angles exceeding $15^{\circ}$ is often insufficient for successful pallet loading. However, as the sensor approaches the pallet, the detection quality improves, maintaining errors within tolerance limits in most cases and enabling a successful interaction.
Figure 10: Pallet pose estimation errors over the distance to sensor $d$ for various load types. Columns represent different orientations around the pallet’s z-axis. Dashed lines indicate tolerance boundaries for successful fork insertion, with the highlighted $0^{\circ}$-$15^{\circ}$ range being critical for insertion.

### v2

@@ 3.2 Actuation
As shown in Table 1, the BM154 platform uses a diesel engine as its main power source, driving two distinct hydraulic circuits that control vehicle movement and operational functions. The diesel engine drives a closed hydraulic circuit designed for operating the wheels and an open hydraulic circuit responsible for actuating multiple hydraulic cylinders, amongst others, for steering and lifting the forks. Figure 4 and Table 2 provide graphical and concise textual information on the forklift’s actuated and non-actuated joints, driven by the hydraulic motor and cylinders.
The closed hydraulic circuit drives the vehicle’s wheels, delivering continuous traction for efficient power transfer and enhanced mobility. This design optimizes vehicle stability and maneuverability, particularly in challenging or uneven terrains. The hydrostatic four-wheel drive system is controlled by a PWM signal, where the duty cycle determines the flow rate of the hydraulic fluid, which in turn directly governs the wheel speed.
@@ 3.3.1 Proprioception
The forklift’s proprioception system integrates several sensors to monitor its internal state with high precision, especially the joint values depicted in Figure 4. In terms of vehicle motion feedback, wheel speed encoders are used to monitor the rotational velocity of each wheel. A system utilizing three linear potentiometers captures the 3D rotation of the central articulated joint. By measuring the sensor elongation, the three articulation angles can be calculated with sub-degree precision. The resulting values are integral for steering the vehicle and precisely adjusting the fork tilt angle during load manipulation. Pressure sensors are installed within the hydraulic cylinders and drive system to measure pressure levels at high frequencies. These measurements are crucial for controlling the forklift’s load-handling capabilities, including detecting whether the fork is in direct contact with the environment. In addition, draw-wire encoders are implemented to measure the vertical position of the lifting mast and the lateral movement of the forks (side-shift) with millimeter precision. Together, these sensors provide comprehensive real-time feedback and allow for accurate modeling of the behavior of the system.
@@ 3.3.2 Exteroception
For pallet recognition, the ZED2i is the central component. It provides the input data for the neural networks for pallet recognition and pose estimation. Additionally, for precise short-range detection during the final pallet approach, the forklift is equipped with a Sick picoScan150 2D LiDAR, Figure. 3d. It ensures accurate positioning of the forks relative to the pallet, due to its high frequent update rate and well-aligned viewing angle. The forklift utilizes the Livox LiDAR to enable accurate detection of the truck’s loading edge. Additionally, a Septentrio mosaic-H system with two antennas is used for RTK GNSS localization. It provides accurate position information at the centimeter level and heading information to ensure precise alignment during loading and unloading tasks.
@@ 3.4 Processing Components and Network
The B&R X90 PLC manages low-level hardware interfaces, including direct connections to sensors and actuators that require low latency and real-time control. It also handles safety-critical operations that depend on rapid response times, such as emergency stops. The rugged IP67-compliant design of the PLC makes it suitable for harsh outdoor environments and does not require an additional enclosure. For higher-level processing, such as planning, advanced control algorithms, and managing intelligent sensors, a Vecow 1210 IPC is employed. It connects to the PLC via a central Ethernet switch to receive the low-level sensor data and to transmit high-level commands over a custom TCP/IP protocol. This setup enables seamless integration and coordination between low-level operations and advanced computational tasks that do not require millisecond-level latency. These advanced computational tasks include high-level task and motion control, vehicle state estimation, and object detection pipelines, utilizing camera and LiDAR sensors. The NUC unit is responsible for processing data from the Ouster LiDAR, allowing both mapping the vehicle’s environment and obstacle avoidance and sharing the respective data via the Robot Operating System (ROS 2) [65] over Ethernet.
A wireless network bridge connects the forklift to a remote monitoring station housed in a weatherproof container, primarily supporting development and expert interventions. This setup enables live monitoring of the system’s operation, ensuring efficient debugging and adjustments during testing. Additionally, a supplementary WLAN antenna allows on-site operators to monitor the system state via a tablet-based Human-Machine Interface (HMI). For long-range communication, an LTE/5G router is integrated.
@@ 3.5 Software Components
Figure 2 shows an overview of the software architecture as well as concise descriptions of core functionalities and their associated hardware components. Core functionalities include localization and mapping, object detection, obstacle detection, task planning, and motion control. Localization and mapping are based on GNSS modules and LiDAR to provide accurate environmental maps and vehicle positioning data. Object detection utilizes a stereo camera and LiDAR to identify pallets and truck loading edges, facilitating precise load carrier handling. The obstacle detection module ensures safety by halting the vehicle before potential collisions with detected obstacles. The task planner serves as the central planning component of the system, managing high-level decision-making and coordinating both the path planner and motion controller. The path planner creates optimal routes based on map and localization data, while the cascaded motion controller provides precise vehicle base movements and responsive fork actuation for effective pallet manipulation. Most software components running on the rugged IPC and the NUC highly depend on the ROS 2 middleware and its associated libraries, which enable a decentralized architecture while ensuring robust data and information exchange. A detailed discussion of the modules presented can be found in Sections 4 and 5.
@@ 4.1 Joint Localization and Pallet Mapping
The goal of mapping is to derive a representation of the environment that is as rich in features as necessary for the application, but as compact as possible to keep computational complexity low. For ADAPT, there are two levels of abstraction in which we need to map the environment: The high-level representation of the pallet poses and the lower-level environment map with a strong focus on terrain traversability assessment, see Section 4.2.
@@
We use a factor graph-based approach [66] that integrates the vehicle’s odometry data, GNSS measurements, and pallet detections into a joint optimization framework for high-level localization and pallet mapping. A factor graph is a theoretical framework that models unknown variables, measurements, and their interdependencies as a bipartite graph with two types of nodes, see Figure 6. It consists of variable nodes (depicted as circles), which represent the variables to be estimated, and factor nodes (depicted as rectangles). In this context, factors can be binary (functions that relate two unknowns with one another) or unary (constraints on a variable based on a measurement). This factor graph approach is advantageous in this context as it allows flexible integration of multiple types of measurement and provides a probabilistic framework for managing uncertainties inherent in noisy and delayed sensor data. Our factor graph is shown in Figure 6. Under typical outdoor conditions, we benefit from the highly accurate GNSS measurements from a dual-antenna system to provide global positioning and attitude information. However, our joint localization and pallet mapping approach enables effective localization of both the vehicle and pallets, even when GNSS signals are degraded or unavailable. While this method remains robust in short-term GPS-denied situations, extended periods without reliable GNSS data will lead to positional drift over time. The forklift pose nodes are connected using odometry based on wheel encoders and steering angle measurements. Pallets are detected and their poses are estimated from different viewpoints according to the forklift poses. Therefore, the estimation of the pose of the pallet is represented as a binary factor linking the position of the pallet with the location of the vehicle. We build upon iSAM2 [33] incremental smoothing to maintain real-time performance, yielding the most probable pallet and forklift poses given all odometry, GNSS, and pallet measurements. This allows the pose of a pallet to be optimized by measurements from different views. In addition, the pallet poses and vehicle location are optimized even without or with a less precise GNSS signal.
@@
With iSAM2 updates of pallet detections and localization information are optimized incrementally, in contrast to repeatedly resolving the entire optimization. To do so, iSAM2 maintains a dynamic Bayesian tree structure to track variable interdependencies in addition to the factor graph. This approach ensures that updates to the factor graph are confined to regions where interdependency changes occur. To facilitate real-time operation, our iSAM2-based optimization approach for joint pallet and vehicle pose mapping
@@
  • exploits the sparsity of the optimization problem, since each factor connects only a small subset of variables, leading to a sparse Jacobian matrix.
Beyond iSAM2 the approach additionally
@@
This approach can efficiently integrate information from odometry sensors, GNSS, and detections over time to yield, given all sensor information, the most probable result for the unknowns (the maximum a posteriori estimate). The pallet mapping is tightly integrated into our optimization framework. For pallet tracking, we use a Mahalanobis distance-based [67] approach to ensure a robust association of detections between consecutive measurements. The Mahalanobis distance metric allows for dynamic association by accounting for sensor pose uncertainty, as well as uncertainty in pallet pose estimation. Additionally, pallets are dynamically managed within the forklift’s view frustum by pallet bookkeeping that goes beyond pure variable marginalization (pruning). It manages the tracking of new never-before-seen pallets, associates new detections with old ones, and deletes pallets if they should be in sensor view but cannot be detected for a while. This approach aligns with the concept of object permanence, where pallets are presumed to remain at a designated location even when they are not directly visible.
@@
Figure 6: Factor Graph for the joint forklift poses ($X_{x}$) and pallet poses ($P_{p}$) over time (left to right), relative (from one time step to the next) odometry information $O_{p1,p2}$, and GNSS measurements $G_{g}$. In this example, the pallet $P_{0}$ was observed in timesteps $1$ and $2$, and GNSS information was available in timesteps 0 and 2.
4.2 Traversability Mapping
@@
The targeted areas of application, e.g., construction sites, of ADAPT are unstructured environments with potentially uneven terrain. Hence, a specific environment mapping targeting these aspects is required. Therefore, a 2.5D elevation map representation is generated from the 3D point measurements of the Ouster LiDAR. While path planning complexity can be reduced to planning in 2 dimensions, the most relevant geometric information about the environment must still be available from the map. In particular, it must allow for a clear distinction between traversable areas, obstacles, and slopes. This distinction is crucial for forklifts, as they are more susceptible to tipping over when carrying loads.
Similarly to spatio-temporal voxel layers [20], our approach aggregates data into a 3D voxel grid with short-term memory. However, instead of directly deriving a 2D cost map, our method employs a 2.5D elevation map for dynamic obstacle handling and terrain analysis, serving as a long-term memory representation. Therefore, it is implemented as a three-stage process, encompassing the analysis of single scans, 3D voxel map aggregation, and the generation of a 2.5D elevation map. In all three steps, distinguishing between ground points and obstacle points plays a decisive role.
Analysis of single scans: Single scans are analyzed within the sensor domain, utilizing the unique capabilities and characteristics of the sensor. By focusing on individual scans, this stage minimizes disturbances caused by alignment errors. Parameters such as the height above ground and the angle between the surface normal vector and the vertical axis help determine the probability of a point being an obstacle. For the Ouster LiDAR, a single swipe of the nearly 180° field of view constitutes a scan. The ego-motion during this rotation is compensated by transforming all points into the sensor’s coordinate system at a fixed timestamp.
@@ 4.3 Pallet Detection and Pose Estimation
Detecting pallets and accurately determining their 6D pose enables ADAPT to interact with them effectively. This is particularly important in construction sites, where pallets are typically placed arbitrarily rather than following a fixed grid or placement schedule.
@@
Pallet representation: Pallets are represented by a part-constellation model, see Figure 9(a), where pallet corners and edges not only form a localized part but at every pixel, they also define a spatial vote offset attribute, implying the offset vector to the corresponding object center. Part locations and local vote vectors are learned solely from synthetic data.
@@
Data generation: A synthetic data generation pipeline is used where more than 150 thousand stereo image pairs are generated with diverse pallet configurations, e.g., orientations, items carried, clutter. The data generation step relies on a simulated stereo camera setup, matching the baseline and optics of a ZED2i stereo camera [68]. The synthetic stereo image pairs are used in a stereo depth computation step, resulting in depth data exhibiting similar characteristics to depth computed from real stereo pairs. This low sim-to-real gap allows for generating a wide variety of view variations and pallet configurations for learning.
Learning: We employ an extensible convolutional encoder-decoder framework, based on [69], which estimates outputs for multiple learning tasks, see Figure 8. Pallets are represented by a part-based representation, where each part also carries a spatial vote offset attribute, where the offset vector points to the corresponding object center. During training, key-point and line segment representations yield multichannel heatmap output representations, where each channel is assigned to respective corner/edge categories.
@@
Pose estimation accuracy: By analyzing the geometry of Euro-pallets and fork dimensions, precise tolerance values for pallet insertion are established. Consequently, pose estimation errors serve as a key metric for evaluating the effectiveness of the proposed approach in achieving accurate insertion. The ground truth data for error computation was obtained through manual annotation of a point cloud captured by an extrinsically calibrated LiDAR sensor mounted adjacent to the camera. Figure 10 presents the evaluation results based on 170 detections, covering various load types at different distances and orientations around the pallet’s z-axis. To compare a detection with its corresponding annotation, the estimated 6D pose was transformed into a coordinate system centered at the manually annotated pose, cf. Figure 9(b), yielding a 6D error vector $\mathbf{e}=[e_{x},e_{y},e_{z},e_{\phi},e_{\theta},e_{\psi}]^{\mathrm{T}}$. The tolerance limits for the relevant degrees of freedom are $y_{tol}=\pm 0.05m$ and $z_{tol}=\pm 0.04m$, with respect to the pallet coordinate system. The constraint in the x-direction is theoretically determined by the load’s center of mass, however, complete fork insertion is preferable to ensure optimal transport stability. Assuming the pallets remain parallel to the ground, detections with unsuitable roll and pitch angles are filtered out, as the forklift cannot successfully pick them up. Thus, these dimensions are excluded from the evaluation. Errors in the y-direction $e_{y}$ are affected by deviations in orientation around the pallet’s z-axis, which modify the effective pallet opening by a factor of $\cos(e_{\psi})$. To account for this influence, the corrected y-direction error is computed as $e_{y}^{\psi}=\frac{e_{y}}{\cos(e_{\psi})}$.
The results indicate that detection accuracy deteriorates with increasing sensor distance, particularly in $e_{x}$, due to the quadratic increase in depth error inherent in stereo vision. Furthermore, when the pallet is rotated around its z-axis, the error distributes between $e_{x}$ and $e_{y}$. Overall, the results suggest that a single detection at angles exceeding $15^{\circ}$ is often insufficient for successful pallet loading. However, the detection quality improves as the sensor approaches the pallet, maintaining errors within tolerance limits in most cases and enabling a successful interaction.
Figure 10: Pallet pose estimation errors over the distance to sensor $d$ for various load types. Columns represent different orientations around the pallet’s z-axis. Dashed lines indicate tolerance boundaries for successful fork insertion, with the highlighted $\psi\in[0^{\circ},15^{\circ}]$ range being critical for insertion.